# Are Large-scale Soft Labels Necessary for Large-scale Dataset Distillation?

### v1

@@ 3.3.1 Diverse Sample Generation via Class-wise Supervision
Since the running statistics are computed per class, we provide the theoretical number of updates required to stabilize all class statistics. We derive the bound using Multiplicative Chernoff Bound and analysis on the momentum decay of statistical significance (see Appendix A for the proof):
$n\geq\max\left(\underbrace{\frac{-2\ln(T)}{\delta^{2}\min\left(p_{c}\right)}}_{\text{from Chernoff}},\underbrace{\frac{\ln(T)}{(1-\delta)\ln(1-\varepsilon)\cdot\operatorname{min}(p_{c})}}_{\text{from decay significance}}\right),$ (8)
where $n$ is the number of BN statistics updates needed, and $p_{c}$ is the probability of class $c$ being sampled from all classes. $T$ is a probability threshold, $\varepsilon$ is the momentum, and $\delta$ is a small positive value, where $0\leq\delta\leq 1$. How Eq. 8 guides our experiment design is detailed in Appendix E.3.
@@ Appendix A Proof
Given that the ImageNet dataset has a varying number of images of classes, we define $p_{c}$ as the probability of sampling from class $c$, where:
$p_{c}=\frac{\text{ number of images in class }c}{\text{ total number of imagesin the dataset }}.$ (9)
We then define $m$ as the number of classes of batches out of a total of $n$ updates (batches) in which class $c$ is included at least once. It follows the Binomial distribution of $n$ trails with a success rate of $p_{c}$:
$m\sim\operatorname{Binomial}\left(n,p_{c}\right).$ (10)
The update rule becomes:
$BN_{c}^{t+m}=(1-\varepsilon)^{m}\cdot BN_{c}^{t}$ (11)
Multiplicative Chernoff Bound. The general bound follows:
$\operatorname{Pr}(X\leq(1-\delta)\mu)\leq e^{-\delta^{2}\mu/2},\quad 0<\delta<1,$ (12)
where $X$ is the sum of independent random variables $X_{i},\dots,X_{n}$ of values 0, 1, and $\mu=E[X]$ denotes the sum’s expected value.
In our specific case, $m$ (the number of batches out of $n$ in which class $c$ is included at least once) suggests that $m$ is the sum of $n$ independent Bernoulli trials, where each trial (or batch) results in a success (1) if class $c$ is present and a failure (0) otherwise. For a small deviation from the expected value, the equation is re-written as:
$\operatorname{Pr}(m\leq(1-\delta)E[m])\leq\exp\left(-\frac{\delta^{2}E[m]}{2}\right)$ (13)
Additionally, we introduce a probability bound that ensures the probability of $m\leq(1-\delta)E[m]$ is bounded by a threshold probability $T$:
$\operatorname{Pr}(m\leq(1-\delta)E[m])\leq\exp\left(-\frac{\delta^{2}E[m]}{2}\right)\leq T$ (14)
Subsitituting $E[m]=n\cdot p_{c}$, we can have:
$\exp\left(-\frac{\delta^{2}\left(n\cdot p_{c}\right)}{2}\right)\leq T$ (15)
To solve for $n$, we take natural log on both sides:
$-\frac{\delta^{2}\left(n\cdot p_{c}\right)}{2}\leq\ln(T)$ (16)
For a small $T$ (i.e., $T<1$), we can rewrite the equation as:
$\frac{\delta^{2}\left(n\cdot p_{c}\right)}{2}$ $\geq-\ln(T)$ (17)
$n\cdot p_{c}$ $\geq\frac{-2\ln(T)}{\delta^{2}}$
$n$ $\geq\frac{-2\ln(T)}{\delta^{2}p_{c}}$
To find the lower bound, we take the probability of the least probability $\operatorname{min}(p_{c})$,
$n\geq\frac{-2\ln(T)}{\delta^{2}\operatorname{min}(p_{c})}$ (18)
Significance Decay. Given $m$ is expected to be at least $(1-\delta)E[m]=(1-\delta)n\cdot p_{c}$ with high probability, we integrate the effects of momentum on the significance of the statistics:
$(1-\varepsilon)^{(1-\delta)n\cdot p_{c}}\cdot BN_{c}^{t}$ (19)
Similarity, we use $T$ as the significance threshold:
$(1-\varepsilon)^{(1-\delta)n\cdot p_{c}}<T$ (20)
$\begin{gathered}(1-\delta)n\cdot p_{c}\cdot\ln(1-\varepsilon)<\ln(T)\\ n\cdot p_{c}>\frac{\ln(T)}{(1-T)\ln(1-\varepsilon)}\\ n>\frac{\ln(T)}{(1-\delta)\ln(1-\varepsilon)\cdot\operatorname{min}(p_{c})}\end{gathered}$ (21)
Combining the two bounds, we have:
@@
$n\geq\max\left(\underbrace{\frac{-2\ln(T)}{\delta^{2}\min\left(p_{c}\right)}}_{\text{from Chernoff}},\underbrace{\frac{\ln(T)}{(1-\delta)\ln(1-\varepsilon)\cdot\operatorname{min}(p_{c})}}_{\text{from decay significance}}\right)$ (22)
@@ E.3 Theoretical Analysis on the Number of Updates
Our experiments are grounded in a careful analysis of the number of updates required for stable Batch Normalization (BN) statistics. We begin by examining Eq. 8:
$n\geq\max\left(\frac{-2\ln(T)}{\delta^{2}\min\left(p_{c}\right)},\frac{\ln(T)}{(1-\delta)\ln(1-\varepsilon)\cdot\min\left(p_{c}\right)}\right)$
To evaluate this equation, we substitute the following values: $T=0.05$ (95% confidence), $\delta=0.2$ (moderate sigma to handle class imbalance), $\epsilon=0.1$ (default value in BN), and $p_{c}=732/1,281,167\approx 0.0005711$ (ratio of the least number of images in a class to total images). We then compute the two parts of the equation separately. The first part, representing the Multiplicative Chernoff Bound, yields:
$\frac{-2\ln(T)}{\delta^{2}\min(p_{c})}=\frac{-2\times\ln(0.05)}{0.2^{2}\times 0.0005711}\approx 262,234$ (23)
The second part, representing the momentum decay of statistical significance, results in:
$\frac{\ln(T)}{(1-\delta)\ln(1-\epsilon)\cdot\min(p_{c})}=\frac{\ln(0.05)}{(1-0.2)\ln(1-0.1)\cdot 0.0005711}\approx 62,234$ (24)
Consequently, we obtain:
$n\geq\max(262,234,62,234)=262,234$ (25)
This theoretical result guides our experiment design in several ways. Firstly, it indicates that the theoretical number of updates needed for stable BN statistics is 262,234. In comparison, the standard ResNet [36] training on ImageNet-1K uses a batch size of 256 and trains for 90 epochs. This results in a total of $(1,281,167/256)\times 90=450,411$ updates, significantly exceeding our theoretical requirement. This observation led to a key insight: pretrained models have already undergone sufficient updates to achieve stable BN statistics. Consequently, we can recompute class-wise BN statistics using a pretrained model for only one epoch (5,005 updates), which is substantially more efficient than retraining from scratch.

### v2

@@ 3.3.1 Diverse Sample Generation via Class-wise Supervision
Since the running statistics are computed per class, we provide the theoretical number of updates required to stabilize all class statistics (see Appendix A for the proof):
$n\geq\max\left(\underbrace{\frac{-2\ln\left(\dfrac{T}{2}\right)}{\delta^{2}\min(q_{c})}}_{\text{Chernoff Bound}},\quad\underbrace{\frac{\ln\left(\dfrac{C}{\tau}\right)}{(1-\delta)\varepsilon\min(q_{c})}}_{\text{BN Convergence}}\right),$ (8)
where $n$ is the number updates needed, $q_{c}$ is the probability that class $c$ appears in a batch, $T$ is a probability threshold, $\varepsilon$ is the momentum parameter in Batch Normalization, $\delta$ is the acceptable relative deviation (where $0\leq\delta\leq 1$), $C$ is some constant, and $\tau$ is the desired convergence tolerance for the BN statistics. How Eq. 8 guides our experiment design is detailed in Appendix E.3.
@@ Appendix A Proof
We aim to determine a lower bound on the number of batches (updates) $n$ required to ensure that the Batch Normalization (BN) statistics for each class in the ImageNet dataset converge within a specified tolerance $\tau$, with high probability. The dataset has a varying number of images per class, affecting the probability of each class appearing in a batch during sampling.
A.1 Preliminary Analysis
Defining Class Probabilities: Let $p_{c}$ denote the probability that a randomly selected image from the dataset belongs to class $c$:
$p_{c}=\frac{\text{Number of images in class }c}{\text{Total number of images in the dataset}}.$
Due to the unequal distribution of images across classes, $p_{c}$ varies among classes.
Probability of Class Appearance in a Batch: When sampling a batch of size $B$, the probability that class $c$ does not appear in the batch is $(1-p_{c})^{B}$. Therefore, the probability that class $c$ appears in the batch is:
$q_{c}=1-(1-p_{c})^{B}.$
This represents the likelihood that at least one image from class $c$ is included in a given batch.
Number of Batches: Let $n$ be the total number of batches sampled during training.
We assume that batches are sampled independently with replacement from the dataset. Under this assumption, each batch is an independent trial where class $c$ appears with probability $q_{c}$. Therefore, the number of batches $M$ where class $c$ appears follows a binomial distribution:
$M\sim\operatorname{Binomial}(n,q_{c}).$
Remark: In practice, batches are often sampled without replacement within an epoch, introducing dependency between batches. However, for large datasets where the total number of images $N$ is significantly larger than the batch size $B$ and the number of batches $n$, the dependence becomes negligible. In such cases, the binomial distribution serves as a reasonable approximation.
The expected value of $M$ is:
$E[M]=nq_{c}.$
A.2 Chernoff Bound
To ensure that $M$ is not significantly less than its expected value $E[M]$, we apply the Chernoff bound:
$\operatorname{Pr}\left(M\leq(1-\delta)E[M]\right)\leq\exp\left(-\frac{\delta^{2}E[M]}{2}\right),$
where $\delta\in(0,1)$ represents the acceptable relative deviation from the expected value. This bound provides a way to quantify the probability that a random variable deviates from its expected value, which is crucial for making high-confidence guarantees.
To ensure the probability that $M$ is less than $(1-\delta)E[M]$ is at most $T_{1}$, we set:
$\exp\left(-\frac{\delta^{2}nq_{c}}{2}\right)\leq T_{1}.$
Solving for $n$:
$n\geq\frac{-2\ln(T_{1})}{\delta^{2}q_{c}}.$
To ensure this condition holds for all classes, we use the minimum value of $q_{c}$:
$n\geq\frac{-2\ln(T_{1})}{\delta^{2}\min(q_{c})}.$
A.3 BN Convergence
BN Statistics Update: Batch Normalization updates its running statistics using an exponential moving average. The update rule for the BN statistics of class $c$ at iteration $t+1$ is:
$\text{BN}_{c}^{t+1}=(1-\varepsilon)\text{BN}_{c}^{t}+\varepsilon\hat{\text{BN}}_{c}^{t+1},$
where $\varepsilon$ is the momentum parameter, and $\hat{\text{BN}}_{c}^{t+1}$ is the BN statistics estimated from the current batch for class $c$.
Since BN statistics for class $c$ are updated only when class $c$ appears in the batch, we consider only the updates corresponding to those batches. After $M$ such updates:
$\text{BN}_{c}^{t+M}=(1-\varepsilon)^{M}\text{BN}_{c}^{t}+\varepsilon\sum_{k=1}^{M}(1-\varepsilon)^{M-k}\hat{\text{BN}}_{c}^{k},$
where $\hat{\text{BN}}_{c}^{k}$ is the BN statistics estimated in the $k$-th batch containing class $c$.
Assuming $\hat{\text{BN}}_{c}^{k}$ is an unbiased estimator of the true BN statistics $\mu_{c}$ when class $c$ appears, the expected value is:
$E[\text{BN}_{c}^{t+M}]=(1-\varepsilon)^{M}\text{BN}_{c}^{t}+\mu_{c}\left[1-(1-\varepsilon)^{M}\right].$
Convergence Within Tolerance: To ensure that the BN statistics converge within a tolerance $\tau$ of the true statistics $\mu_{c}$:
$\left|E[\text{BN}_{c}^{t+M}]-\mu_{c}\right|\leq\tau.$
Since:
$\left|E[\text{BN}_{c}^{t+M}]-\mu_{c}\right|=(1-\varepsilon)^{M}\left|\text{BN}_{c}^{t}-\mu_{c}\right|,$
and assuming $\left|\text{BN}_{c}^{t}-\mu_{c}\right|\leq C$ for some constant $C$, we have:
$(1-\varepsilon)^{M}C\leq\tau.$
Taking natural logarithms:
$M\ln(1-\varepsilon)+\ln(C)\leq\ln(\tau).$
Using the approximation $\ln(1-\varepsilon)\approx-\varepsilon$ for small $\varepsilon$:
$-M\varepsilon+\ln(C)\leq\ln(\tau).$
Solving for $M$:
$M\geq M_{0}=\frac{\ln\left(\dfrac{C}{\tau}\right)}{\varepsilon}.$
Origin of $M_{0}$: Here, $M_{0}$ is derived from the BN convergence requirement that ensures:
$(1-\varepsilon)^{M_{0}}\left|\text{BN}_{c}^{t}-\mu_{c}\right|\leq\tau.$
It represents the minimum number of updates required for the BN statistics of class $c$ to converge within the desired tolerance $\tau$.
A.4 Combining Bounds
Event Definitions:
  • Let $E_{1}$ be the event that class $c$ appears in sufficient batches (as guaranteed by the Chernoff bound).
  • Let $E_{2}$ be the event that the BN statistics for class $c$ converge within the desired tolerance $\tau$.
Target Probability: We aim to ensure that both events occur simultaneously with high probability:
$P(E_{1}\cap E_{2})\geq 1-T.$
Union Bound Application: For any two events, the probability of their intersection satisfies:
$P(E_{1}\cap E_{2})$ $=1-P(\overline{E_{1}\cap E_{2}})$
$=1-P(\overline{E_{1}}\cup\overline{E_{2}})$
$\geq 1-P(\overline{E_{1}})-P(\overline{E_{2}}).$
Error Probability Allocation: For simplicity, we allocate the total acceptable failure probability $T$ equally between the two events:
[…]
Chernoff Bound Analysis: For event $E_{1}$, we require that the probability of class $c$ appearing in fewer than the expected number of batches is at most $\frac{T}{2}$:
$P\left(M\leq(1-\delta)nq_{c}\right)\leq\frac{T}{2}.$
Applying the Chernoff bound:
$\exp\left(-\frac{\delta^{2}nq_{c}}{2}\right)\leq\frac{T}{2}.$
Solving for $n$:
$-\frac{\delta^{2}nq_{c}}{2}$ $\leq\ln\left(\frac{T}{2}\right),$
$n$ $\geq\frac{-2\ln\left(\frac{T}{2}\right)}{\delta^{2}q_{c}}.$
BN Convergence Requirement: For event $E_{2}$, we require that the number of batches $M$ where class $c$ appears is sufficient for BN convergence:
$M\geq M_{0}=\frac{\ln\left(\dfrac{C}{\tau}\right)}{\varepsilon}.$
To ensure that this condition holds when event $E_{1}$ occurs, we use the fact that, with probability at least $1-\frac{T}{2}$, we have:
$M\geq(1-\delta)nq_{c}.$
Therefore, to guarantee $M\geq M_{0}$, we require:
$(1-\delta)nq_{c}\geq M_{0}=\frac{\ln\left(\dfrac{C}{\tau}\right)}{\varepsilon}.$
Solving for $n$:
$n$ $\geq\frac{\ln\left(\dfrac{C}{\tau}\right)}{(1-\delta)\varepsilon q_{c}}.$
Final Combined Bound: To ensure that both conditions hold for all classes, we use $\min(q_{c})$:
@@ A.4 Combining Bounds
$n\geq\max\left(\underbrace{\frac{-2\ln\left(\dfrac{T}{2}\right)}{\delta^{2}\min(q_{c})}}_{\text{Chernoff Bound}},\quad\underbrace{\frac{\ln\left(\dfrac{C}{\tau}\right)}{(1-\delta)\varepsilon\min(q_{c})}}_{\text{BN Convergence}}\right),$
where $\delta$ represents the acceptable relative deviation from the expected number of batches, $\varepsilon$ is the momentum parameter in BN updates, $T$ denotes the acceptable total failure probability ($T=T_{1}+T_{2}$), $\tau$ is the convergence threshold for BN statistics, $C$ represents an upper bound on $\left|\text{BN}_{c}^{t}-\mu_{c}\right|$ at initialization, and $\min(q_{c})$ represents the minimum probability that a class appears in a batch.
This bound ensures:
  • With probability at least $1-\frac{T}{2}$, each class $c$ appears in at least $(1-\delta)nq_{c}$ batches (event $E_{1}$ occurs).
  • With probability at least $1-\frac{T}{2}$, the BN statistics for each class $c$ converge within tolerance $\tau$ (event $E_{2}$ occurs).
  • By the union bound, both events $E_{1}$ and $E_{2}$ occur simultaneously with probability at least $1-T$.
@@ E.3 Theoretical Analysis on the Number of Updates
Our experiments are grounded in a careful analysis of the number of updates required for stable Batch Normalization (BN) statistics. We begin by examining the derived bound from Eq. 8:
$n\geq\max\left(\underbrace{\frac{-2\ln\left(\dfrac{T}{2}\right)}{\delta^{2}\min(q_{c})}}_{\text{Chernoff Bound}},\quad\underbrace{\frac{\ln\left(\dfrac{C}{\tau}\right)}{(1-\delta)\varepsilon\min(q_{c})}}_{\text{BN Convergence}}\right).$
To evaluate this bound, we substitute the following values:
  • $T=0.05$ (acceptable total failure probability, corresponding to 95% confidence)
  • $\delta=0.2$ (acceptable relative deviation from the expected number of batches)
  • $\varepsilon=0.1$ (momentum parameter in BN)
  • $\min(p_{c})=\dfrac{732}{1,281,167}\approx 0.0005711$ (ratio of the least number of images in a class to total images)
  • $B=256$ (batch size)
  • $\min(q_{c})=1-(1-\min(p_{c}))^{B}$ (minimum probability that any class appears in a batch)
First, we compute $\min(q_{c})$:
$\min(q_{c})$ $=1-(1-\min(p_{c}))^{B}$
$=1-\left(1-0.0005711\right)^{256}$
$=1-\left(0.9994289\right)^{256}$
$\approx 1-e^{-256\times 0.0005711}\quad\text{(since $\min(p_{c})$is small)}$)
$=1-e^{-0.1462}$
$\approx 1-0.8639=0.1361.$
Thus, $\min(q_{c})\approx 0.1361$.
Next, we compute the two parts of the bound separately.
From Chernoff Bound Term: Given that we allocate the total failure probability $T$ equally between the two events, we have $T/2=0.025$.
$n$ $\geq\frac{-2\ln\left(\dfrac{T}{2}\right)}{\delta^{2}\min(q_{c})}$
$=\frac{-2\ln(0.025)}{(0.2)^{2}\times 0.1361}$
$=\frac{-2\times(-3.6889)}{0.04\times 0.1361}\quad\text{(since $\ln(0.025)=-3.6889$)}$
$=\frac{7.3778}{0.005444}$
$\approx 1,355.2.$
From BN Convergence Term: We need to specify $C$ and $\tau$. Let’s assume:
  • $C=1$ (an upper bound on $\left|\text{BN}_{c}^{t}-\mu_{c}\right|$ at initialization, as the running mean is typically initialized to zero)
  • $\tau=0.01$ (desired convergence tolerance)
Compute the numerator:
$\ln\left(\dfrac{C}{\tau}\right)=\ln\left(\dfrac{1}{0.01}\right)=\ln(100)=4.6052.$
Now, compute the denominator:
$(1-\delta)\varepsilon\min(q_{c})=(1-0.2)\times 0.1\times 0.1361=0.8\times 0.1\times 0.1361=0.010888.$
Compute the second part:
$n$ $\geq\frac{\ln\left(\dfrac{C}{\tau}\right)}{(1-\delta)\varepsilon\min(q_{c})}$
$=\frac{4.6052}{0.010888}$
$\approx 423.08.$
Final Bound:
[…]
This theoretical result indicates that approximately $1,356$ batches are needed for stable BN statistics with the specified parameters.
Practical Implications: This observation leads to a key insight: pretrained models have already undergone sufficient updates to achieve stable BN statistics. Specifically, in the context of ImageNet-1K:
$\text{Updates per epoch}=\frac{1,281,167}{256}\approx 5,005\,\text{updates}>1,356.$
Since one epoch consists of approximately $5,005$ updates, which is substantially more than the theoretical requirement of $1,356$ batches, we can confirm that a single epoch of training is sufficient for the BN statistics of each class to converge within the desired tolerance with high probability.